OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (1, 1).

OpenStudy (bahrom7893):

again use what we did 2ce so far: y-y1 = M(x-x1)

OpenStudy (anonymous):

iknow how im just lazy

OpenStudy (bahrom7893):

don't be lazy.. come on.

OpenStudy (anonymous):

igot no paper blah

OpenStudy (bahrom7893):

neither do I.. Come one just plug in points and slope and u're done!

OpenStudy (bahrom7893):

Ughh okay here: y-1 = -3(x-1) y-1 = -3x+3 y = -3x + 3 +1 y = -3x+4

OpenStudy (bahrom7893):

was that hard?

OpenStudy (anonymous):

nope, thankks